OpenStudy (anonymous):

Write the given expression in terms of x and y only. cos(arcsin x − arctan y)

OpenStudy (freckles):

let a=arcsin(x) and b=arctan(y)

OpenStudy (freckles):

cos(a-b) equals what in terms of cos(a),cos(b),sin(a), and sin(b)

OpenStudy (freckles):

use the difference identity for cosine that is

OpenStudy (freckles):

now for sin(a)=x and tan(b)=y we are going to build two separate right triangles

OpenStudy (freckles):

sin(a)=x/1 so for this triangle you will label the side opp of angle a x and hyp 1 ----- tan(b)=y/1 so for this triangle you will label the side opp of angle b y and the adjacent side of angle b 1

OpenStudy (freckles):

you should find the other unknown side in both triangles

OpenStudy (freckles):

you will use this info write cos(a-b) as an algebraic expression

OpenStudy (anonymous):

my answer didnt work for that

OpenStudy (freckles):

have you expanded cos(a-b) using the difference identity?

OpenStudy (anonymous):

yes unless i did the math wrong, but i used cosacosb+sinasinb and tried to simplify after that

OpenStudy (freckles):

\[\cos(a-b)=\cos(a)\cos(b)+\sin(a)\sin(b)\] that is correct and we have this also